OpenStudy (anonymous):

Simplify ln(lne^e^10)

OpenStudy (liamschumm):

ln means log base e.

geerky42 (geerky42):

\[\ln\left(\ln\left(e^{e^{10}}\right)\right)\]?

OpenStudy (anonymous):

Yes, minus the parentheses around e^e^10, but I don't think that really matters. Then again, I suck at math so it could.

OpenStudy (aum):

Use the fact that \(\ln(e^x) = x\) \[ \ln\left(\ln e^{e^{10}}\right) = \ln\left(e^{10}\right) = 10 \]